# Zero shot Generalization of Vision-Based RL Without Data Augmentation

### v1

@@ A.4 Beta Study
Small values of $\beta$ result in a more even distribution of the probability mass between latent values per codebook, which implies that the output will be a weighted sum of different latent values (or memories under the Hopfield interpretation). We choose three different values, $\beta=$ (1, 10, 50), and compare with the main result ($\beta=100$) presented in the paper on the Walker domain, shown in 12. To our surprise, lower $\beta$ values have little to no effect on generalization performance and, in fact, increase training performance. This perhaps challenges the assumptions made in Hsu et al. (2023) about the requirements of disentanglement via latent quantization, but admittedly requires further analysis, which we leave to future work.
Figure 12: ALDA is trained on the Walker domain with different values of $\beta$, with evaluations periodically performed on the "color hard" and DistractingCS environments. Average of 4 seeds, shaded region represents a 95% bootstrapped confidence interval.
@@ A.5 Framestack Ablation
We provide an ablation comparing against a version of ALDA where the encoder receives as input, and the decoder predicts a stack $k=3$ of frames, i.e., the observation size is $(9\times 64\times 64)$. Since the downstream 1D-CNN layer is no longer necessary, we remove this layer from the variant. We refer to this variant as "ALDA (framestack)" and present results on the Walker domain in Figure 13. We also provide latent traversal visualizations of ALDA (framestack), shown in Figure 14.
@@
A.6 Hyperparameters and Model Architectures
Our SAC implementation is based on Yarats & Kostrikov (2020).
A.6.1 Actor and Critic Networks
Following Yarats & Kostrikov (2020), we use double $Q$-networks, each of which is a 3-layer multi-layer perceptron (MLP) with 1024 hidden units per hidden layer and GeLU activations after all except the final layer. The actor network is similarly a 3 layer MLP with 1024 hidden units per layer and GeLU activations on all but the final layer.
A.6.2 Encoder, Decoder, and Latent Model
We use the same encoder/decoder architectures as Hsu et al. (2023), with the exception that we replace all leaky ReLU activations with GeLU. We instantiate the codebooks for the latent model with values evenly spaced between [-1, 1].
A.6.3 Hyperparemeters

### v2

@@ A.4 Beta Study
Small values of $\beta$ result in a more even distribution of the probability mass between latent values per codebook, which implies that the output will be a weighted sum of different latent values (or memories under the Hopfield interpretation). We choose three different values, $\beta=$ (1, 10, 50), and compare with the main result ($\beta=100$) presented in the paper on the Walker domain, shown in 12. To our surprise, lower $\beta$ values have little to no effect on generalization performance and, in fact, increase training performance. This perhaps challenges the assumptions made in (Hsu et al., 2023) about the requirements of disentanglement via latent quantization, but admittedly requires further analysis, which we leave to future work.
Figure 12: ALDA is trained on the Walker domain with different values of $\beta$, with evaluations periodically performed on the ”color hard” and DistractingCS environments. Average of 4 seeds, shaded region represents a 95% bootstrapped confidence interval.
@@ A.5 Framestack Ablation
We provide an ablation comparing against a version of ALDA where the encoder receives as input, and the decoder predicts a stack $k=3$ of frames, i.e., the observation size is $(9\times 64\times 64)$. Since the downstream 1D-CNN layer is no longer necessary, we remove this layer from the variant. We refer to this variant as ”ALDA (framestack)” and present results on the Walker domain in Figure 13. We also provide latent traversal visualizations of ALDA (framestack), shown in Figure 14.
@@
A.6 Implementation Details
Our SAC implementation is based on (Yarats & Kostrikov, 2020).
A.7 DMControl Generalization Benchmark
The DMControl Generalization Benchmark (DMCGB) (Hansen & Wang, 2021) is an extension of the DeepMind Control Suite (DMControl) (Tassa et al., 2018). DMControl is a software stack consisting of physics simulation for a suite of common reinforcement learning, continuous control environments. DMCGB introduces visual distribution shifts to benchmark visual generalization performance of models trained on these environments from image observations.
A.7.1 Actor and Critic Networks
Following (Yarats & Kostrikov, 2020), we use double $Q$-networks, each of which is a 3-layer multi-layer perceptron (MLP) with 1024 hidden units per hidden layer and GeLU activations after all except the final layer. The actor network is similarly a 3 layer MLP with 1024 hidden units per layer and GeLU activations on all but the final layer.
A.7.2 Encoder, Decoder, and Latent Model
We use the same encoder/decoder architectures as (Hsu et al., 2023), with the exception that we replace all leaky ReLU activations with GeLU. We instantiate the codebooks for the latent model with values evenly spaced between [-1, 1].
A.7.3 Hyperparemeters
@@
A.8 ALDA Pseudocode
Algorithm 1 ALDA Forward Pass
Input: Observation o, encoder $f_{\theta}$, latent model $l_{\psi}$, history encoder $h_{\gamma}$.
$\textbf{o}\in\mathbb{R}^{B\times Ck\times H\times W}\rightarrow\textbf{o}\in\mathbb{R}^{Bk\times C\times H\times W}$ // rearrange the framestack dimension
$z_{cont}\in\mathbb{R}^{Bf\times n_{z}}\leftarrow f_{\theta}(\textbf{o})$
$z_{d}\in\mathbb{R}^{Bf\times n_{z}}\leftarrow l_{\psi}(z_{cont})$ // association step using the latent model
$z_{d}\in\mathbb{R}^{Bf\times n_{z}}\rightarrow z_{d}\in\mathbb{R}^{B\times f\times n_{z}}$ // rearrange the framestack dimension
$z_{d}\leftarrow h_{\gamma}(z_{d})$ // encode temporal information return $z_{d}$
Algorithm 2 Associative Latent Dynamics
Input: Continuous latent vector $z_{cont}\in\mathbb{R}^{Bk\times n_{z}}$, latent model $l_{\psi}$
for $i\leftarrow 1\;\textbf{to}\;n_{z}$ do
  $\textbf{w}_{i}\leftarrow\text{Softmax}(-\beta L_{1}(z_{cont_{i}},\textbf{v}_{i}))\odot\textbf{v}_{i}$ // compute weights for how similar $z_{cont_{i}}$ is to each value in the $i^{\prime}th$ codebook.
  $z_{d_{i}}\leftarrow\sum_{|\textbf{v}_{i}|}\textbf{w}_{i}$
end forreturn $z_{d}$
Algorithm 1 contains pseudocode for ALDA’s forward pass. The observation o can be an in-distribution sample during training, or an OOD sample during evaluation. Post-training, when presented with in-distribution samples, the association step is unlikely to significantly change $z_{cont}$, since $z_{cont}$ will map very close to the values learned by the latent model $l_{\psi}$. However, when presented with OOD samples, $z_{cont}$ is more likely to change since certain dimensions of $z_{cont}$ may map far away from the corresponding dimensions of the latent model. Algorithm 2 shows how the latent model $l_{\psi}$ maps potentially OOD continuous latent vectors to in-distribution values using modern Hopfield retreival mechanisms.
A.9 Latent Study
Figure 15: Training (left) and evaluation (middle and right) results of increasing the dimensionality of ALDA’s latent space on the Walker Walk task. Results are averaged over 4 seeds. Shaded region represents a 95% bootstrapped confidence interval.
We examine the effects of increasing the dimensionality of the latent space on the performance of the Walker ”Walk” task and present the results in Figure 15. The baseline model ($z_{d}=12$) performs the best on the training and ”Color Hard” evaluation tasks, and that performance drops as the dimensionality of the latent space increases. Since disentangled representation learning methods try to approximate the number of ground truth sources of variation of the data distribution, it is possible that setting $z_{d}$ to values far away from the number of true sources can cause the performance degradation we observe in Figure 15.
A.10 Critic Gradient Ablation
Figure 16: Ablation of backpropogating the critic gradients to the encoder (ALDA CG) compared to the standard model. Results are averaged over 4 seeds. Shaded area represents a 95% bootstrapped confidence interval.
In this experiment, we examine the effects of backpropagating the critic’s gradients through the latent model and back to the encoder. The results are presented in Figure 16. We refer to the ALDA variant with critic gradients enabled as ”ALDA (CG)” and compare on the Walker ”Walk” task. ALDA (CG) performs worse on all training and evaluation tasks. We suspect that backpropagating the critic gradients to the encoder affects the ability for the encoder-decoder pair to disentangle sources of the image distribution, since disentangled representation learning methods typically study disentanglement in (Variational) Autoencoders without competing auxiliary objectives.
A.11 SVEA with Other Data Augmentations
Figure 17: Comparison of ALDA to SVEA (image overlay) and SVEA (color easy), where SVEA (color easy) is trained directly on the color easy environment as a data augmentation.
SVEA (image overlay) leverages the Places dataset of 1 million realistic images, which likely puts the DMCGB evaluation tasks in-distribution. However, how well a model handles OOD data i.e. how well it performs extrapolative generalization is an important consideration for generalist agents. In this study, we train SVEA directly on the color easy evaluation environment in DMCGB as a form of weaker data augmentation (DA). Since the color easy environment only slightly randomizes the scene RGB values, it puts the color hard and Distracting CS environments out-of-distribution for SVEA, and allows us to tease out how well data augmentation helps with extrapolative generalization. The results are presented in Figure 17.
We find that SVEA (color easy) performs worse than both SVEA (image overlay) and ALDA on the training and Color Hard environments, and equals ALDA in performance on Distracting CS. This suggests that not all forms of DA are equally useful, and that DA should be used with caution e.g. when the range of possible variations on the data is bounded and it is computationally tractable to expose the model to all variations. While DA has shown to be useful on toy benchmarks, datasets, or in constrained lab settings, the results presented here suggest that this approach will not scale to, for example, in-the-wild deployments in the real world where the data scale and variance are practically unbounded. The results also highlight the need for models that are robust to distribution shifts without knowing or training on the distribution shifts apriori.